# CoVariance Filters and Neural Networks

### v1

@@ 1 Introduction
Contribution. We introduce a novel convolutional learning framework for random signals defined over general, possibly infinite-dimensional, Hilbert spaces, centered on the covariance operator of such space. We leverage the eigensystem of the covariance operator as a frequency domain to define a Hilbert coVariance Fourier Transform, and then we propose spectral Hilbert coVariance Filters (HVF) to manipulate the signal representation in such spectral domain. We show how one can seamlessly develop Hilbert coVariance Networks (HVNs) as a layered architecture of HVFs and nonlinear activations in the same spirit as GNNs [10]. We propose a principled discretization procedure to make HVFs and HVNs implementable, and we prove that empirical HVFs can recover the Functional PCA (FPCA) [22] of the filtered signals. Finally, we numerically validate HVNs on both synthetic and real-world binary classification tasks of time-series, corroborating our findings and showing robustness benefits w.r.t. FPCA and covariance-unaware learning such as MLPs.
@@ 3 Hilbert Covariance Filters and Neural Networks
This operation, where the frequency response $h(\lambda_{\ell})$ amplifies or attenuates the different functional principal components, is in complete analogy with the convolutions/spectral filtering in several other domains, e.g., time [25], graphs [26], or manifolds [27]. By learning or designing $h(\lambda_{\ell})$, we are able to put different weights on the different HVFT coefficients. As such, taking only the $r$ variance-maximizing directions, i.e., employing the FPCA transform as a dimensionality-reduction method, is then equivalent to an ideal high/low-pass HVF (if we order the eigenvalues of the covariance operator in ascending/descending order). General spectral filters are maximally expressive, but they require explicit computation of eigenfunctions and eigenvalues. For this reason, spatial (polynomial) filters are often preferred in practice, as they are (i) scalable and stable–they avoid eigendecomposition and are eigenvector-agnostic, and (ii) sill expressive–polynomials have strong approximation capabilities.
@@
A HVN with input signals $\left\{x^{i}\right\}_{i=1}^{F_{0}}$ is then built as the stack of $T$ layers defined as in (7), where $x_{0}^{i}=x^{i}$. To globally represent the HVN, we collect all the filters in $\mathcal{W}=\left\{h_{t}^{u,i}\right\}_{t,u,i}$ and describe the HVN outputs as a mapping $\{x_{T}^{u}\}_{u=1}^{F_{T}}=\bm{\Psi}\left(\mathcal{W},C,\left\{x^{i}\right\}_{i=1}^{F_{0}}\right)$ to enhance that it is parameterized by filters $\mathcal{H}$ and the covariance operator $C$. An HVN can be spectral or spatial, depending on the type of employed filters.
@@ 4 Discretization via Bounded Operators
Proof. See Appendix.
Discrete HVFs, and HVNs. Once signals are discretized and the empirical covariance matrix is computed, discrete HVFs and HVNs are simply given by (6) and (7) with the covariance operator replaced by the empirical covariance matrix and the signals replaced by their discretized version, i.e. $\{x^{(m),u}_{T}\}_{u=1}^{F_{T}}=\bm{\Psi}(\mathcal{W},\hat{\mathbf{C}}_{n}^{(m)},\{x^{(m)}_{i}\}_{i=1}^{n})$. Discretized HVNs are principled instances of coVariance Neural Network (VNN) [9] and, as a consequence, of Graph (convolutional) Neural Networks (GNNs) [10]. To better enhance this similarity, we rewrite the layer in (7) in matrix form by introducing the matrices $\mathbf{X}^{(m)}_{t}=\{x_{t,u}^{(m)}\}_{u=1}^{F_{t}}\in\mathbb{R}^{m\times F_{l}}$, and $\mathbf{W}_{t,j}=\{w_{t,j}^{u,q}\}_{q=1,u=1}^{F_{l},F_{l+1}}\in\mathbb{R}^{F_{l}\times F_{l+1}}$, as
$\mathbf{X}^{(m)}_{t+1}=\sigma\Bigg{(}\sum_{j=0}^{J}\Big{(}\hat{\mathbf{C}}_{n}^{(m)}\Big{)}^{j}\mathbf{X}_{t}^{(m)}\mathbf{W}_{t,j}\Bigg{)}\;\in\mathbb{R}^{m\times F_{l+1}},$ (17)
@@
Use-Cases for (Discrete) HVNs. To showcase the versatility of our framework, in Appendix B we illustrate three practical $(\mathcal{H},S)$ pairs together with admissible nonlinear activations $\sigma:\mathcal{H}\to\mathcal{H}$. In particular, we describe how HVNs can be applied to functional data (e.g. time series, as we also show in Sec. 5), infinite sequences, and data defined over Reproducing Kernel Hilbert Spaces.
@@ 5 Numerical Results
$\textrm{and }\qquad\Sigma_{c}^{(0)}=I_{d},\qquad\Sigma_{c}^{(1)}=\big{[}\rho^{|i-j|}\big{]}_{i,j=1}^{d},$ (19)
@@
where $\phi,\rho>0$. In this task, a data point with class $y$ consists of a bag of $n$ i.i.d. samples drawn from $x_{1},\dots,x_{n}\ \stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\ \mathsf{GP}\big{(}0,K^{(y)}\big{)},$ discretized as noisy $x_{i}^{(m)}:=S_{m}x_{i}+\epsilon\in\mathbb{R}^{m}$, where $\epsilon$ is AWGN, and centered by $\overline{x}^{(m)}=\tfrac{1}{n}\sum_{i=1}^{n}x_{i}^{(m)}$. We compare HVNs against two baselines acting on the same discretized bags: (i) MLP, obtained by simply setting $J=1$ and $\hat{\mathbf{C}}_{n}^{(m)}=I_{m}$ in (7), and (ii) FPCA, consisting of computing the first $O$ discrete FPCA scores (i.e., the PCA scores of the discrete $\{x_{i}^{(m)}\}_{n}$ in the bag). For all the tested models, the output of the last layer (for HVN and MLP) or the scores (for PCA) are fed to an MLP classifier to get the class probabilities. We set $\phi=0.20$, and $\rho=0.7$, train with ADAM optimizer to minimize cross–entropy loss, we normalize the compressed covariance matrices for training stability (for HVN), and we use GELU nonlinearity (for HVN and MLP), mean pooling, and a polynomial degree (for HVN) $J=2$. For both HVN and MLP, we use 2 layers, 32 output signals in each HVN layer, and we adjust the MLP hidden dimensions to match the number of parameters of HVN. We fix $d=4$ and $p=32$, and we generate balanced train/test sets for different numbers of samples $n$ and signal to noise ratios $SNR_{dB}$, defined, given a bag, as $10\log_{10}\left((1/n)\sum_{i=1}^{n}\|x_{i}^{(m)}\|^{2}/\textrm{var}(\epsilon)\right)$. In Figures 1(a)-1(b) we show the test accuracy vs $n$ (for fixed $SNR_{dB}=30$) and vs $SNR_{dB}$ (for fixed $n=24$), respectively. The MLP achieves mediocre accuracies (and is prone to overfitting) as it never mixes the $m$ components of the samples, accessing only first–order information, whereas the label is encoded in cross–feature covariance. The FPCA baseline gives random accuracy as it rotates each bag to its variance–maximizing basis, leading to coordinate drift across samples (sign flips, permutations,…), which a single classifier must reconcile. In contrast, HVN shows strong performance thanks to the transferable nature of HVFs and their ability to capture discriminative second–order structure consistently across bags. We further evaluate HVNs on the ECG5000 dataset from the UCR repository [28], which contains 5000 ECG samples of length 140 and 5 classes denoting the heartbeat type. We again compare HVNs, MLP, and FPCA with bin averaging and varying discretization resolution $m$. The hyperparameters are set as in the synthetic experiment. In this case, the goal is to classify individual time-series and not (possibly streaming) bags; thus, both for HVNN and FPCA, we compute the empirical covariance matrix on the entire training (and test, at inference) set and then use it to classify the individual time-series. In this way, both HVNN and FPCA have a more robust estimate of the covariance operator. Indeed, Figure 1(c) shows that HVN and FPCA outperform MLP across all values of $m$, indicating the importance of covariance information. HVN outperforms both MLP and FPCA by a margin, especially for increasing $m$, corroborating its capability to incorporate covariance information into its end-to-end learning dynamics robustly.

### v2

@@ 1 Introduction
Contribution. We introduce a novel convolutional learning framework for random signals defined over general, possibly infinite-dimensional, Hilbert spaces, centered on the covariance operator of such space. We leverage the eigensystem of the covariance operator as a frequency domain to define a Hilbert coVariance Fourier Transform, and then we propose spectral Hilbert coVariance Filters (HVF) to manipulate the signal representation in such spectral domain. We show how one can seamlessly develop Hilbert coVariance Networks (HVNs) as a layered architecture of HVFs and nonlinear activations in the same spirit as GNNs [10]. We propose a principled discretization procedure to make HVFs and HVNs implementable, and we prove that empirical HVFs can recover the Functional PCA (FPCA) [22] of the filtered signals. Finally, we numerically validate HVNs on both synthetic and real-world classification tasks of time-series, corroborating our findings and showing robustness benefits w.r.t. FPCA and covariance-unaware learning such as MLPs.
@@ 3 Hilbert Covariance Filters and Neural Networks
This operation, where the frequency response $h(\lambda_{\ell})$ amplifies or attenuates the different functional principal components, is in complete analogy with the convolutions/spectral filtering in several other domains, e.g., time [25], graphs [26], or manifolds [27]. By learning or designing $h(\lambda_{\ell})$, we are able to put different weights on the different HVFT coefficients. As such, taking only the $r$ variance-maximizing directions, i.e., employing the FPCA transform as a dimensionality-reduction method, is then equivalent to an ideal high pass HVF. General spectral filters are maximally expressive, but they require explicit computation of eigenfunctions and eigenvalues. For this reason, spatial (polynomial) filters are often preferred in practice, as they are (i) scalable and stable–they avoid eigendecomposition and are eigenvector-agnostic, and (ii) still expressive–polynomials are strong approximators.
@@
A HVN with input signals $\left\{x^{i}\right\}_{i=1}^{F_{0}}$ is then built as the stack of $T$ layers defined as in (7), where $x_{0}^{i}=x^{i}$. To globally represent the HVN, we collect all the filters in $\mathcal{W}=\left\{h_{t}^{u,i}\right\}_{t,u,i}$ and describe the HVN outputs as a mapping $\{x_{T}^{u}\}_{u=1}^{F_{T}}=\bm{\Psi}\left(\mathcal{W},C,\left\{x^{i}\right\}_{i=1}^{F_{0}}\right)$ to enhance that it is parameterized by filters $\mathcal{H}$ and the covariance operator $C$. An HVN can be spectral or spatial, depending on the type of filters employed.
@@ 4 Discretization via Bounded Operators
Proof. See Appendix A.
Discrete HVFs, and HVNs. Once signals are discretized and the empirical covariance matrix is computed, discrete HVFs and HVNs are simply given by (6) and (7) with the covariance operator replaced by the empirical covariance matrix and the signals replaced by their discretized version, i.e. $\{x^{(m),u}_{T}\}_{u=1}^{F_{T}}=\bm{\Psi}(\mathcal{W},\hat{\mathbf{C}}_{n}^{(m)},\{x^{(m)}_{i}\}_{i=1}^{n})$. Discretized HVNs are principled instances of coVariance Neural Network (VNN) [9] and, as a consequence, of Graph (convolutional) Neural Networks (GNNs) [10]. To better enhance this similarity, we rewrite the layer in (7) in matrix form by introducing the matrices $\mathbf{X}^{(m)}_{t}=\{x_{t,i}^{(m)}\}_{i=1}^{F_{t}}\in\mathbb{R}^{m\times F_{t}}$, and $\mathbf{W}_{t,j}=\{w_{t,j}^{i,q}\}_{q=1,i=1}^{F_{l},F_{t+1}}\in\mathbb{R}^{F_{l}\times F_{t+1}}$, as
$\mathbf{X}^{(m)}_{t+1}=\sigma\Bigg(\sum_{j=0}^{J}\Big(\hat{\mathbf{C}}_{n}^{(m)}\Big)^{j}\mathbf{X}_{t}^{(m)}\mathbf{W}_{t,j}\Bigg)\;\in\mathbb{R}^{m\times F_{t+1}},$ (17)
@@
Use-Cases for (Discrete) HVNs. To showcase the versatility of our framework, in Appendix B, we illustrate three practical $(\mathcal{H},S)$ pairs together with admissible nonlinear activations $\sigma:\mathcal{H}\to\mathcal{H}$. In particular, we describe how HVNs can be applied to functional data (e.g. time series, as we also show in Sec. 5), infinite sequences, and data defined over Reproducing Kernel Hilbert Spaces.
@@ 5 Numerical Results
$\textrm{and }\hskip 18.49988pt\Sigma_{c}^{(0)}=I_{d},\hskip 18.49988pt\Sigma_{c}^{(1)}=\big[\rho^{|i-j|}\big]_{i,j=1}^{d},$ (19)
@@
where $\phi,\rho>0$. In this task, a data point with class $y$ consists of a bag of $n$ i.i.d. samples drawn from $x_{1},\dots,x_{n}\ \stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\ \mathsf{GP}\big(0,K^{(y)}\big),$ discretized as noisy $x_{i}^{(m)}:=S_{m}x_{i}+\epsilon\in\mathbb{R}^{m}$, where $\epsilon$ is AWGN, and centered by $\overline{x}^{(m)}=\tfrac{1}{n}\sum_{i=1}^{n}x_{i}^{(m)}$. We compare HVNs against two baselines acting on the same discretized bags: (i) MLP, obtained by simply setting $J=1$ and $\hat{\mathbf{C}}_{n}^{(m)}=I_{m}$ in (7), and (ii) FPCA, consisting of computing the first $O$ discrete FPCA scores (i.e., the PCA scores of the discrete $\{x_{i}^{(m)}\}_{n}$ in the bag). For all the tested models, the output of the last layer (for HVN and MLP) or the scores (for PCA) are fed to an MLP classifier to get the class probabilities. We set $\phi=0.20$, and $\rho=0.7$, train with ADAM optimizer to minimize cross–entropy loss, we normalize the compressed covariance matrices for training stability (for HVN), and we use GELU nonlinearity (for HVN and MLP), mean pooling, and a polynomial degree (for HVN) $J=2$. For both HVN and MLP, we use 2 layers, 32 output signals in each HVN layer, and we adjust the MLP hidden dimensions to match the number of parameters of HVN. We fix $d=4$ and $p=32$, and we generate balanced train/test sets for different numbers of samples $n$ and signal to noise ratios $SNR_{dB}$, defined, given a bag, as $10\log_{10}\left((1/n)\sum_{i=1}^{n}\|x_{i}^{(m)}\|^{2}/\textrm{var}(\epsilon)\right)$. In Figures 1(a)-1(b) we show the test accuracy vs $n$ (for fixed $SNR_{dB}=30$) and vs $SNR_{dB}$ (for fixed $n=24$), respectively. The MLP achieves mediocre accuracies (and is prone to overfitting) as it never mixes the $m$ components of the samples, accessing only first–order information, whereas the label is encoded in cross–feature covariance. The FPCA baseline gives random accuracy as it rotates each bag to its variance–maximizing basis, leading to coordinate drift across samples (sign flips, permutations,…), which a single classifier must reconcile. In contrast, HVN shows strong performance thanks to the transferable nature of HVFs and their ability to capture discriminative second–order structure consistently across bags. We further evaluate HVNs on the ECG5000 dataset from the UCR repository [28], which contains 5000 ECG samples of length 140 and 5 classes denoting the heartbeat type. We again compare HVNs, MLP, and FPCA with bin averaging and varying discretization resolution $m$. The hyperparameters are set as in the synthetic experiment. In this case, the goal is to classify individual time-series and not (possibly streaming) bags; thus, both for HVNN and FPCA, we compute the empirical covariance matrix on the entire training set and then use it to classify the individual time-series. In this way, both HVNN and FPCA have a more robust estimate of the covariance operator. Indeed, Figure 1(c) shows that HVN and FPCA outperform MLP across all values of $m$, indicating the importance of covariance information. Moreover, HVN outperforms both MLP and FPCA by a margin, especially for increasing $m$, corroborating its capability to robustly incorporate covariance information into its end-to-end learning dynamics.